Suppose the price elasticity of demand for heating oil is 0.2 in the short run and 0.7 in the long run.

a. Ifthepriceofheatingoilrisesfrom$1.80to$2.20

per gallon, what happens to the quantity of heating oil demanded in the short run? In the long run? (Use the midpoint method in your calculations.) 


1
Expert's answer
2022-02-14T15:32:06-0500

"∆In Quantity=Price Elasticity×(\frac{∆InPrice}{2})"

Therefore in the short run,

"∆InQuantity=0.2×(\frac{2.2-1.8}{2})=0.04"

In the Long run,

"∆InQuantity=0.7×(\frac{2.2-1.8}{2})=0.14"